OpenStudy (anonymous):

the symbol\[A _{b}\] stands for the projection of vector A onto vector B. In other words, \[A _{b}\] represents the component of A that is parallel to B. Derive an expression for \[A _{b}\] in terms of the vectors A and B.

OpenStudy (amistre64):

\[\frac{a.b}{|a|^2}a\]

OpenStudy (anonymous):

wow u did that really quickly

OpenStudy (amistre64):

just got out of calc3 that just went over it :)

OpenStudy (anonymous):

in the numerator...is that a times b?

OpenStudy (amistre64):

a dot b

OpenStudy (anonymous):

a dot b divided by magnitude of a squared all multiplied by a?

OpenStudy (amistre64):

yes, the left side is a scalar; and the right side is vector a scaled to that length

OpenStudy (anonymous):

would it be a bother to run me through how u got to this expression?

OpenStudy (amistre64):



OpenStudy (amistre64):

|b| cos(t) = the length of Ab, if we use your notation for proj{a} b

OpenStudy (amistre64):

\[|b|cos(t) =|b| \frac{a.b}{|a||b|}=\frac{a.b}{|a|}\] good so far?

OpenStudy (anonymous):

yea i get that

OpenStudy (anonymous):

i'm with u

OpenStudy (anonymous):

i'm also trying to decode your picture :)

OpenStudy (anonymous):

is that the x coordinate down there at the end?

OpenStudy (anonymous):

the long vector

OpenStudy (amistre64):

now, we need to scale that to a unit vector of a; since a is |a| long, lets divide it by |a| to get it to a length of 1

OpenStudy (amistre64):

\[\frac{a.b}{|a|}\ \frac{a}{|a|}=; unit\ a, scaled\ by\ needed\ length\]

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

and since |a| |a| = |a|^2 i just condensed it alittle bit

OpenStudy (anonymous):

simple enough?

OpenStudy (amistre64):